OpenStudy (anonymous):

Solve EXACTLY in terms of log. 10/(1+10^-x) = 2 x= ?

OpenStudy (anonymous):

\[\frac{10}{1+10^{-x}} =2\]\[\implies 10 = 2(1+10^{-x})\] Can you solve from here?

OpenStudy (anonymous):

10= 2(1+1/10x) ?

OpenStudy (anonymous):

That's a good start, though it's 10^x not 10x

OpenStudy (anonymous):

been a while so thanks for the patience

OpenStudy (anonymous):

ah so 1/10^x

OpenStudy (anonymous):

So: \[10 = 2(1+\frac{1}{10^x})\]

OpenStudy (anonymous):

Now distribute the 2 to each term in parens.

OpenStudy (anonymous):

10= 2+2/10^x ?

OpenStudy (anonymous):

so 10= 2 + 1/8^x ?